OpenStudy (fantalux):

Someone plz help! 38-2x=6

OpenStudy (kainui):

Any ideas? Give it your best try and we'll help you.

OpenStudy (anonymous):

38-2x = 6 38-2x-38 = 6 - 38 -2x = -32 x = 16

OpenStudy (fantalux):

I added 2 to six like I was supposed to do, but this is a multiple choice question and my answer that I came out with wasn't there.

OpenStudy (fantalux):

The both of you.

OpenStudy (fantalux):

Thx